OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 The third step would be to: Select one: a. Graph the function. b. Take the log of each side. c. Square the expression. d. Factor the resulting quadratic equation.

OpenStudy (anonymous):

@phi @kc_kennylau

OpenStudy (kc_kennylau):

log_2(x) + log_2(x-7) = 3 log_2[x(x-7)] = 3 x(x-7) = 8 x^2-7x = 8 <----- third step x^2-7x-8 = 0 (x-8)(x+1) = 0 x=8 or x=-1(rej)

OpenStudy (kc_kennylau):

There are so many approaches, I don't understand why particularly ask for the third step...

OpenStudy (anonymous):

is it take the log of each side?

OpenStudy (kc_kennylau):

@satellite73

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@satellite73